OpenStudy (anonymous):

write an equation for the parabola whose vertex is at (3,3) and passes through (5,27)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

y = a(x-h)^2 +k vertex: (h,k) = (3,3) y=a(x-3)^2 +3 use (5,27) for x and y, solve for a 27 = a(5-3)^2 +3 27 = 4a +3 24 = 4a 6 = a y = 6(x-3)^2 + 3

myininaya (myininaya):

good job cow! :)